A bag contians 4 black marbles, 6 red marbles, 9 green marbles, and 5 blue marbles. What is the probability of pulling 2 consecutive red marbles?
F. 5/92
G. 7/92
H. 9/92
J. 11/92

Answer by checkley77(12569)

6/24*5/23=30/552=5/92 ANSWER F).